OpenStudy (anonymous):

Arrange the following growth rates in increasing order: (2 marks) O ( (24)n ) , O (n2 ), O (1), O (n log n)

OpenStudy (anonymous):

Lol, test questions, really?

OpenStudy (anonymous):

So you have \[ O(24n), O(n^2), O(1), O(n\log n) \]

OpenStudy (anonymous):

First of all \(O(1)\) is the lowest it gets.

OpenStudy (anonymous):

To compare the other ones divide by \(n\) to get: \(24, n, \log n\) In this case it is clear that \(24\) is next lowest leaving you with \(n\) and \(\log n\).

OpenStudy (anonymous):

It's fairly obvious that \(\log n\) grows slower than \(n\).